# Complex-Valued Holographic Radiance Fields

### v1

@@ 4.3. Geometrically Aligned Phase Modeling
where $R(x,y)$ is the reference wave. While a properly estimated 3D hologram inherently contains information for different viewing angles within the limits of the pixel pitch— functioning as a “window” into a virtual world that observers can view from different positions— the issue arises when the geometric relationship between the object and the hologram plane changes. Unlike assuming a stationary or narrowly perturbed observer at the center of projection, any significant movement of the camera plane itself within the 3D world demands recalculating the hologram:
@@
where $\theta$ represents new geometric configuration when relationship between scene geometry and viewpoint changes. Existing methods estimate, optimize, or calculate closed-form solutions of $P(x,y)$ to produce desired intensity patterns, treating each geometric configuration as an independent optimization problem. As shown in  Fig. 2, when the scene-to-hologram relationship changes, the previously computed hologram leads to visuals that are highly distorted, resembling noise. Although neural networks enable efficient recalculation of holograms for new configurations, these learned methods still fail to maintain geometrically aligned phase across different camera and geometric arrangements. Instead, neural networks essentially create a new, unrelated data-driven solution, which lacks interpretability and physical consistency; in real-world scenarios, the phase relationships within the same object remain invariant regardless of its geometric relationship to the observation plane.
Figure 4. Comparison of different hologram synthesis methods across different geometric arrangements in simulation. The top row shows our method maintaining geometrically aligned phase from large left to right rotations for novel views. The middle rows show the traditional methods that rely on intensity-based intermediaries, which fail to maintain physical consistency. The bottom row shows the intensity ground truth. * We reimplement the methodology of (Choi et al., 2025) to demonstrate the rotation of single-view propagation results; original non-rotational results show higher image quality than ours.
In traditional intensity-based rasterization, each pixel depends only on a handful of nearby fragments; in , by contrast, every primitive contributes across a large “sub-hologram” footprint on the full resolution hologram, and one must accumulate complex-valued wavefronts from all primitives at all spatial frequencies. Primitive-based methods that maintain physical accuracy (e.g.,  (2025) or ploygon-based ) require proper handling of occlusions and interference effects through complex wavefront accumulation:
@@
While this reduces complexity to $O(N_{\text{primitives}}\times N_{\text{freq}})$ and enables better parallelization, it sacrifices accurate occlusion modeling and still requires substantial computation, as each primitive must be re-evaluated across the entire hologram plane. Existing methods follow an Eulerian computational paradigm, where the hologram plane coordinates remain fixed while complex fields are recalculated for each viewpoint.
@@
Geometrically Aligned Phase.
In contrast, our approach fundamentally differs by modeling the complex field directly in 3D space using Gaussian primitives with intrinsic phase properties — a Lagrangian perspective where the content remains fixed while the geometric relationship changes. Each complex-valued Gaussian $\mathcal{G}_{n}$ has an explicitly defined phase parameter $\boldsymbol{\varphi}_{n}$ that represents its inherent holographic properties, independent to camera parameters. The complex field contribution of a Gaussian remains consistent with respect to its 3D position, orientation, and intrinsic phase. This Lagrangian formulation enables us to avoid recalculations in the Eulerian approach, abandons the expensive per-frequency summation, and instead pushes all of the work into two highly optimized GPU-friendly stages: 1. Tile-based Complex-Valued Rasterization: As described in Sec. 4.2, by dividing the screen into 16×16 tiles, we extend the rasterizer to accumulate complex-valued Gaussians directly in image space and reduce the rasterization cost to $O(N_{\text{primitives}})$. 2. FFT-based layer propagation: Instead of summing each primitive’s contribution at every spatial frequency, we collect all Gaussians onto $L$ depth planes once, and then propagate these planes to the hologram plane using 2D FFTs. This reduces the propagation cost to $O(L\times N_{\text{res}}\log N_{\text{res}})$, where $N_{\text{res}}=n_{x}\times n_{y}$ is the number of pixels per depth plane, independent to the number of primitives. In this case, since $L$ does not grow with $N_{\text{primitives}}$ and $L\times N_{\text{res}}\log N_{\text{res}}$ is much smaller than $N_{\text{primitives}}\times N_{\text{freq}}$ when the number of primitives is large, our approach achieves inherently better scaling. This results in a more straightforward holographic rendering pipeline that maintains constant propagation cost and geomterically aligned phase regardless of scene complexity. Mathematically, consider two different geometric configurations with respective projection matrices $\mathbf{J}_{1}$, $\mathbf{W}_{1}$ and $\mathbf{J}_{2}$, $\mathbf{W}_{2}$. For a given $\mathcal{G}_{n}$, while its projected complex field will differ between geometric configurations, the underlying phase relationship remains consistent:
(24) $U_{n}^{(1)}$ $=\mathbf{c}_{n}\mathcal{G}_{n}^{\text{proj}}(\mathbf{J}_{1},\mathbf{W}_{1})\exp(j\boldsymbol{\varphi}_{n})$
@@
where $U_{\Pi_{l}}^{(i)}$ represents the complex field at depth plane $\Pi_{l}$ for geometric configuration $i$, and $P^{(i)}$ is the recorded hologram. As shown in  Fig. 4, the significance of the view-consistent phase becomes apparent when examining the reconstructed wavefront across different viewpoints. In real life, light emitted or scattered from a physical point maintains a geometrically aligned phase relationship relative to other points, regardless of the geometric arrangement or observer’s position. Complex-valued holographic radiance field attempts to mimic this physical behavior, ensuring that the phase differences between scene elements remain invariant, leading to geometrically aligned interference and diffraction effects. This represents a paradigm shift—from treating the phase as a computational convenience applied after intensity determination on the hologram plane, to modeling it as an intrinsic and meaningful property of the scene itself. By combining the Lagrangian perspective of 3DGS with holography, we achieve modeling and rendering of a complex-valued holographic radiance field, eliminating the need for expensive per-configuration, per-primitive wave calculations.
@@ 5. Implementation
We employ the ground truth images and rendered depth information to generate target reconstruction images using the multiplane generation pipeline by Kavaklı  et al. (2023b). Additionally, we employ the fused SSIM proposed by (Mallick et al., 2024) to speed up the training process. For densification, we adopt the original strategy from , with the exceptions that the densifying frequency is 300 steps and the opacity resetting is replaced by the regularization method proposed by (Rota Bulò et al., 2024).
We choose Adan (Xie et al., 2024) as the optimizer ($\beta_{1}=0.9,\beta_{2}=0.99$) and train our complex-valued Gaussians for 20000 steps using parameter-specific learning rates: $0.005$ for scales, $0.0025$ for both amplitude and phase, $0.025$ for opacity, $0.001$ for rotation, and a base learning rate of $0.01$ for means and plane assignment probabilities. Our choice of optimizer follows (Zhang et al., 2024), while the learning rate is adopted from the original implementation. We apply CosineAnnealingLR (2017) only to the learning rates of means and plane assignment probabilities (minimum $0.00001$) for smoother convergence, while keeping the others fixed throughout training. For holographic configuration, we follow the most recent literature (Shi et al., 2022, 2021; Akşit and Itoh, 2023) and choose a propagation distance of 2 $mm$, volume depth of 4 $mm$, pixel pitch of 3.74 $\mu m$, and wavelengths of 639, 532, and 473 $nm$.
Table 2. Performance comparison on NeRF Synthetic and LLFF. GT represents the 3DGS + U-Net (GT), which is Image Quality Ground Truth that requires recalculation per viewpoint. 3DGS + U-Net (Vary) represents the Viewpoint Relationship Variation without recalculation. Metrics are reported as PSNR/SSIM/LPIPS, respectively. Bold indicates best performance excluding GT baseline. For more information about baselines, please refer to Sec. 6.1.
@@ 6. Evaluation
We evaluate our method on three standard datasets to demonstrate its effectiveness, including NeRF Synthetic (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022).
Table 3. Comparison of inference times across different Gaussian primitive based hologram synthesis methods at 800×800 resolution.
@@ 6.1. Baseline Selection
We compare our method against a two-stage method that generates 3D holograms using a U-Net (Ronneberger et al., 2015) from RGBD inputs queried from the pretrained intensity-based 3DGS scene. This two-stage method serves as our baseline for several reasons. First, while neural network-based hologram estimation lacks geometrically aligned phase across viewpoints, it achieves millisecond-level inference time that is substantially faster than other methods, such as optimization. Second, although the learned method introduces unnatural defocus effects and suffers from a limited eye box, the network’s denoising capacity helps minimize speckle noise and artifacts commonly found in other CGH methods, resulting in the top metric image quality. This represents the theoretical upper bound of image quality that our method can achieve under ideal conditions. Therefore, we use the 3DGS + U-Net (GT) method as our benchmark, serving as the Image Quality Ground Truth against which the performance of our method is evaluated. Additionally, we evaluate the Viewpoint Relationship Variation scenario where 3DGS + U-Net generates holograms without recalculation when the relationship between scene geometry and viewpoint changes, named 3DGS + U-Net (Vary).
@@ 6.2. Inference Time
Table. 3 demonstrates our method’s efficiency. Compared with , our method is inherently more scalable, achieving 30x-10,000x speedup while preserving geometric consistency across novel views. While 3DGS + U-Net offers faster rendering, it requires a re-estimation of the phase per viewpoint, compromising geometric consistency. Even with a large number of Gaussians (5M), our method maintains reasonable performance (69 $ms$), achieving a balance between computational efficiency and the representation of a view-consistent complex-valued holographic radiance field. Fig. 5 illustrates the scalability of our method; the render time remains relatively constant across resolutions due to tile-based rasterization, while both memory usage and render time scale linearly with the number of depth planes and Gaussian primitives, demonstrating the computational efficiency of our method compared to traditional methods that rely on per-primitive diffraction calculation.
@@ 6.3. Complex Field Discontinuities: Beyond Intensity-Based Smooth Interpolation
6.3.1. Quantitaive Analysis.
Table. 2 and 4 show the training images performance of our method and 3DGS + U-Net on the NeRF Synthetic, LLFF, and Mip-NeRF 360 datasets, respectively. We evaluate the image quality using PSNR, SSIM, and LPIPS (Zhang et al., 2018a).
The Viewpoint Relationship Variation results show that 3DGS + U-Net (Vary) fails to generate correct holograms when geometry changes without recalculation, with PSNR dropping by 21.5 dB, 20.4 dB, and 19.6 dB on NeRF Synthetic, LLFF, and Mip-NeRF 360 datasets, respectively, and SSIM decreases by over 0.74 across all 17 scenes. In contrast, our method maintains robust performance under geometric variations, achieving 18.7 dB, 18.5 dB, and 13.9 dB PSNR improvements over 3DGS + U-Net (Vary).
For the Image Quality Ground Truth, our method achieves promising performance on the NeRF Synthetic and LLFF datasets, with PSNR gaps of 2.8 dB and 1.9 dB and SSIM differences of 0.05 and 0.11, respectively, compared to 3DGS + U-Net (GT). However, performance degradation of our method becomes significant on Mip-NeRF 360 datasets with 5.7 dB PSNR drop and 0.42 SSIM decrease. This progressive degradation from controlled synthetic scenes to complex in-the-wild environments stems from the fundamental challenge of representing spatially discontinuous complex fields to model interference and diffraction.
@@
In intensity-based radiance fields, the goal is to model view-dependent appearance through opacity and color. This representation handles sparse viewpoint sampling effectively because intensity variations generally follow smooth, low-frequency patterns that can be interpolated between viewpoints. However, complex-valued holographic radiance fields modeling introduces a challenge related to motion parallax. For a camera translation, motion parallax causes objects at different depths to experience different amounts of apparent displacement. Objects at depths $d_{1}$ and $d_{2}$ exhibit displacements of:
@@
where $f$ is the focal length and $\Delta t$ is the camera translation. When $d_{2}\gg d_{1}$, the motion parallax difference $\Delta s_{1}-\Delta s_{2}$ becomes significant. This motion parallax translates directly into phase differences in complex-valued holographic radiance fields. Between consecutive camera viewpoints, these primitives accumulate different phase shifts:
@@
When the scene contains large depth discontinuities ($d_{2}\gg d_{1}$), the spatial phase gradient will become very large. This creates interpolation challenges absent in intensity-based representations. Unlike intensity values, which are continuous and monotonic, phase values wrap at $2\pi$ boundaries. Large spatial phase gradients cause adjacent pixels to differ by multiple cycles, creating ambiguity in interpolation paths that intensity representations don’t encounter. To evaluate our method against complex field discontinuities, we carefully choose and analyze challenging scenes with significant depth variations from commonly used datasets that exhibit different motion parallax characteristics.
@@
As shown in Fig. 6, the NeRF Synthetic dataset benefits from controlled capture conditions with relatively uniform depth distributions within each scene. In contrast, LLFF and Mip-NeRF 360 datasets contain in-the-wild scenes with larger depth variations. These scenes create sharp phase boundaries where the complex field changes abruptly in between viewpoints, breaking the spatial coherence assumptions of intensity-based representation. While the 3DGS + U-Net method maintains visual quality by first modeling geometry and appearance through intensity-based 3DGS (which handles motion parallax effectively) and then estimating phase as a secondary task, this two-stage approach sacrifices the physical consistency across different geometric configurations. This limitation highlights the trade-off between the spatial coherence requirements of intensity-based representations and the phase discontinuities introduced by motion parallax in holographic radiance fields.
Table 4. Performance comparison on Mip-NeRF 360. GT means the 3DGS + U-Net (GT), which is Image Quality Ground Truth that requires recalculation per viewpoint. 3DGS + U-Net (Vary) means the Viewpoint Relationship Variation without recalculation. Metrics are reported as PSNR/SSIM/LPIPS, respectively. Bold indicates best performance excluding GT baseline.

### v2

@@ 4.3. Scene Geometry-Aware Amplitude and Phase Representations
where $R(x,y)$ is the reference wave. While a properly estimated 3D hologram inherently contains information for different viewing angles within the Space-Bandwidth Product (SBP) limits of the pixel pitch— functioning as a “window” into a virtual world that observers can view from different positions— the issue arises when the geometric relationship between the scene geometry and the hologram plane changes. Unlike assuming a stationary or narrowly perturbed observer at the center of projection, any significant movement of the camera plane itself within the 3D world demands recalculating the hologram:
@@
where $\theta$ represents the new viewpoint when the geometric relationship between scene geometry and hologram plane changes. Existing CGH methods estimate, optimize, or calculate closed-form solutions of $P(x,y)$ to produce desired intensity patterns, treating each viewpoint as an independent optimization problem. As shown in  Fig. 2, when the scene-to-hologram relationship changes, the previously computed hologram leads to visuals that are highly distorted, resembling noise. Although neural networks enable efficient recalculation of holograms for new configurations, these learned methods fail to maintain consistent amplitude and phase across different viewpoints. Instead, neural networks essentially create a new, unrelated data-driven solution, which lacks interpretability and physical consistency; in real-world scenarios, the amplitude and phase relationships from an object remain invariant regardless of its geometric relationship to the observation plane.
Figure 4. Comparison of different hologram synthesis methods across different viewpoints in simulation. The top row shows our method with scene geometry-aware representations from large left to right rotations for novel views. The middle rows show the existing CGH methods that rely on intensity-based intermediaries, which fail to maintain consistency across novel views. The bottom row shows the intensity ground truth. * We reimplement (Choi et al., 2025) to demonstrate the results of GWS; The original non-rotational result shows higher image quality and different defocus blur than our reimplementation.
In traditional intensity-based rasterization, each pixel depends only on a handful of nearby fragments; in CGH, by contrast, every primitive contributes across a large “sub-hologram” footprint on the full resolution hologram, and one must accumulate complex-valued wavefronts from all primitives at all spatial frequencies. Primitive-based methods that maintain physical accuracy (e.g., GWS (Choi et al., 2025) or polygon-based CGH) require proper handling of occlusions and interference effects through complex wavefront accumulation:
@@
While this reduces complexity to $O(N_{\text{primitives}}\times N_{\text{freq}})$ and enables better parallelization, it sacrifices accurate occlusion modeling and still requires substantial computation, as each primitive must be re-evaluated across the entire hologram plane. Existing CGH methods follow an Eulerian computational paradigm, where the hologram plane coordinates remain fixed while complex fields are recalculated for each viewpoint.
@@
Our Method.
In contrast, our approach differs fundamentally by modeling the complex field directly in 3D space using Gaussian primitives with intrinsic phase properties — a Lagrangian perspective where the content remains fixed while the viewpoint changes. Each complex-valued Gaussian $\mathcal{G}_{n}$ has an explicitly defined phase parameter $\boldsymbol{\varphi}_{n}$ that represents its inherent holographic properties, independent to camera parameters. We emphasize that this per-Gaussian phase parameter does not correspond to the absolute optical phase measured at the sensor in real life and does not change with the position of the observer or the camera. Instead, it serves as a learned intrinsic reference phase, characterizing the local wave emission of each Gaussian within the scene‘s Lagrangian coordinate. View-dependent and position-dependent interference will arise subsequently through light propagation. Although the intrinsic phase parameters remain fixed across viewpoints, the result of interference observed at the camera plane will vary consistently with scene geometry. Therefore, the complex field contribution of each Gaussian is uniquely determined by its 3D position, orientation, and intrinsic phase, remaining coherent across views.
This Lagrangian formulation enables us to avoid recalculations in the Eulerian approach, avoids the expensive per-frequency summation, and instead pushes all of the work into two highly optimized GPU-friendly stages: 1. Tile-based Complex-Valued Rasterization: As described in Sec. 4.2, by dividing the screen into 16×16 tiles, we extend the 3DGS rasterizer to accumulate complex-valued Gaussians directly in image space and reduce the rasterization cost to $O(N_{\text{primitives}})$. 2. FFT-based layer propagation: Instead of summing each primitive’s contribution at every spatial frequency, we collect all Gaussians onto $L$ depth planes once, and then propagate these planes to the hologram plane using 2D FFTs. This reduces the propagation cost to $O(L\times N_{\text{res}}\log N_{\text{res}})$, where $N_{\text{res}}=n_{x}\times n_{y}$ is the number of pixels per depth plane, independent to the number of primitives. In this case, since $L$ does not grow with $N_{\text{primitives}}$ and $L\times N\log N$ is much smaller than $N_{\text{primitives}}\times N_{\text{freq}}$ when the number of primitives is large, our approach achieves inherently better scaling. This results in a more straightforward holographic rendering pipeline that maintains constant propagation cost regardless of scene complexity. Mathematically, consider two different viewpoints with respective projection matrices $\mathbf{J}_{1}$, $\mathbf{W}_{1}$ and $\mathbf{J}_{2}$, $\mathbf{W}_{2}$. For a given $\mathcal{G}_{n}$, while its projected complex field will differ between viewpoints, the underlying phase relationship remains consistent:
(24) $U_{n}^{(1)}$ $=\mathbf{c}_{n}\mathcal{G}_{n}^{\text{proj}}(\mathbf{J}_{1},\mathbf{W}_{1})\exp(j\boldsymbol{\varphi}_{n}),$
@@
where $U_{\Pi_{l}}^{(i)}$ represents the complex field at depth plane $\Pi_{l}$ for viewpoint $i$, and $P^{(i)}$ is the recorded hologram. As shown in  Fig. 4, this consistency becomes apparent when examining the reconstructed wavefront across different viewpoints. In real life, light maintains intrinsic phase offsets at emission that are tied to scene geometry, while the observed phase relationships at the sensor vary with viewpoint through wave propagation and superposition (Jang et al., 2024). Complex-valued holographic radiance field attempts to mimic this physical behavior, ensuring that the amplitude and phase differences between scene elements remain invariant, leading to physically-consistent interference and diffraction effects. This represents a paradigm shift—from treating the amplitude and phase as computational variables applied after intensity determination on the hologram plane, to modeling them as an intrinsic and meaningful property of the 3D scene itself. By combining the Lagrangian perspective of 3DGS with holography, we achieve modeling and rendering of a complex-valued holographic radiance field, eliminating the need for expensive per-configuration, per-primitive wave calculations.
@@ 5. Implementation
We employ the ground-truth images and rendered depth information to generate target reconstruction images using the multiplane generation pipeline by Kavaklı  et al. (2023b). Additionally, we employ the fused SSIM proposed by (Mallick et al., 2024) to speed up the training process. For densification, we adopt the original strategy from 3DGS, with the exceptions that the densifying frequency is 300 steps and the opacity resetting is replaced by the regularization method proposed by Rota Bulò et al. (2024).
We choose Adan (Xie et al., 2024) as the optimizer ($\beta_{1}=0.9,\beta_{2}=0.99$) and train our complex-valued Gaussians for 20000 steps using parameter-specific learning rates: $0.005$ for scales, $0.0025$ for both amplitude and phase, $0.025$ for opacity, $0.001$ for rotation, and a base learning rate of $0.01$ for means and plane assignment probabilities. Our choice of optimizer follows (Zhang et al., 2024), while the learning rate is adopted from the original 3DGS implementation. We apply CosineAnnealingLR (Loshchilov and Hutter, 2017) only to the learning rates of means and plane assignment probabilities (minimum $0.00001$) for smoother convergence, while keeping the others fixed throughout training. For holographic configuration, we follow the most recent literature (Shi et al., 2022, 2021; Akşit and Itoh, 2023) and choose a propagation distance of 2 $mm$, volume depth of 4 $mm$, pixel pitch of 3.74 $\mu m$, and wavelengths of 639, 532, and 473 $nm$. All experiments are conducted on a NVIDIA A100 80G GPU. For details on dataset-specific camera viewpoint sampling strategies, please refer to Supplementary Sec. C.
Table 2. Quality comparison on NeRF Synthetic and LLFF. 3DGS + U-Net (GT) refers to the Image Quality Ground Truth that requires recalculation of hologram per viewpoint. 3DGS + U-Net (Vary) represents the Viewpoint Relationship Variation without recalculation. Metrics are reported as PSNR/SSIM/LPIPS, respectively. Bold indicates best performance excluding GT baseline. For more information about baselines, please refer to Sec. 6.1.
@@ 6. Evaluation
We evaluate our method on three standard datasets, including NeRF Synthetic (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022).
Table 3. Comparison of inference times across different Gaussian primitive based hologram synthesis methods at 800×800 resolution. Inference time for 3DGS + U-Net includes per-view recomputation, consisting of 3DGS rendering and network inference for each queried viewpoint.
@@ 6.1. Baseline Selection
We compare our method against a two-stage baseline that generates holograms using a U-Net (Ronneberger et al., 2015) from RGBD images rendered from a pretrained intensity-based 3DGS scene. The baseline is selected for several reasons. First, although neural network estimates holograms that are not scene-geometric aware, it achieves millisecond-level inference, substantially faster than optimization-based CGH methods. Second, despite introducing unnatural defocus effects and a limited eye box, the network’s denoising capability suppresses speckle noise and artifacts commonly found in the other CGH methods, yielding the highest quantitative image quality. As such, this method represents a theoretical upper bound on achievable image quality under ideal conditions. We therefore treat 3DGS + U-Net (GT) as the Image Quality Ground Truth for benchmarking. In addition, we evaluate a Viewpoint Relationship Variation setting, denoted 3DGS + U-Net (Vary), where holograms are generated without recalculation when the relationship between scene geometry and viewpoint changes.
From a computational perspective, 3DGS + U-Net requires rendering RGBD images from the pretrained 3DGS scene at each queried viewpoint, followed by neural inference to generate the hologram. Similarly, GWS first renders the 2DGS scene, then computes closed-form wave propagation for each Gaussian primitive at the queried viewpoint. Both adopt an Eulerian paradigm, recalculating holograms for every novel view at camera plane. We use 3DGS + U-Net as the primary baseline because the source code of GWS is currently unavailable; our reimplementation yields lower image quality, and its reconstructions exhibit different defocus blur characteristics than the official implementation, making the comparison inaccurate and unfair.
@@ 6.2. Inference Time
Table 3 demonstrates our method’s efficiency. Compared with GWS (Choi et al., 2025), our method is inherently more scalable, achieving 30x-10,000x speedup while maintaining view consistency. While 3DGS + U-Net (Chen et al., 2025) offers faster rendering, it requires recalculation of the hologram per viewpoint, which has no scene geometry-awareness. Even with a large number of Gaussians (5M), our method maintains reasonable performance (69 $ms$), achieving a balance between computational efficiency and maintaining scene geometry-aware representations. Fig. 5 illustrates the scalability of our method; the render time remains relatively constant across resolutions due to tile-based rasterization, while both memory usage and render time scale linearly with the number of depth planes and Gaussian primitives, demonstrating the computational efficiency of our method compared to traditional CGH methods that rely on per-primitive diffraction calculation. For ablation study of the computational contribution of tile-based rasterization and FFT-based propagation, please refer to Supplementary Sec. E.
@@ 6.3. Complex Field Discontinuities: Beyond Intensity-Based Smooth Interpolation
6.3.1. Quantitative Analysis.
Table 2 and 4 show the training images performance of our method and 3DGS + U-Net on the NeRF Synthetic, LLFF, and Mip-NeRF 360 datasets, respectively. We evaluate the image quality using PSNR, SSIM, and LPIPS (2018a).
The Viewpoint Relationship Variation results show that 3DGS + U-Net (Vary) fails to generate correct holograms when viewpoint changes without recalculation, with PSNR dropping by 21.5 dB, 20.4 dB, and 19.6 dB on NeRF Synthetic, LLFF, and Mip-NeRF 360 datasets, respectively, and SSIM decreases by over 0.74 across all 17 scenes. In contrast, our method maintains robust performance across novel views, achieving 18.7 dB, 18.5 dB, and 13.9 dB PSNR improvements over 3DGS + U-Net (Vary).
For the Image Quality Ground Truth, our method achieves promising performance on the NeRF Synthetic and LLFF datasets, with PSNR gaps of 2.8 dB and 1.9 dB and SSIM differences of 0.05 and 0.11, respectively, compared to 3DGS + U-Net (GT). However, performance degradation of our method becomes significant on Mip-NeRF 360 dataset with 5.7 dB PSNR drop and 0.42 SSIM decrease. This progressive degradation from controlled synthetic scenes to complex in-the-wild environments stems from the fundamental challenge of representing spatially discontinuous complex fields to model interference and diffraction. For more analysis of gaussian distribution and statistics of holograms rendered by our method, please refer to Supplementary Sec. F.
@@
In intensity-based radiance fields, the goal is to model view-dependent appearance through opacity and color. This representation handles sparse viewpoint sampling effectively because intensity variations generally follow smooth, low-frequency patterns that can be interpolated between viewpoints. However, modeling complex-valued holographic radiance fields introduce a challenge related to motion parallax. For a camera translation, motion parallax causes objects at different depths to experience different amounts of apparent displacement. Objects at depths $d_{1}$ and $d_{2}$ exhibit displacements of:
@@
where $f$ is the focal length and $\Delta t$ is the camera translation. When $d_{2}\gg d_{1}$, the motion parallax difference $\Delta s_{1}-\Delta s_{2}$ becomes significant. This motion parallax translates directly into phase differences in complex-valued holographic radiance fields. Between consecutive camera viewpoints, these Gaussian primitives will accumulate different phase shifts:
@@
When the scene contains large depth discontinuities ($d_{2}\gg d_{1}$), the spatial phase gradient will become very large. This creates interpolation challenges absent in intensity-based representations. Unlike intensity values, which are continuous and monotonic, phase values wrap at $2\pi$ boundaries. Large spatial phase gradients cause adjacent pixels to differ by multiple cycles, creating ambiguity in interpolation paths that intensity-based representations do not encounter. To evaluate our method against complex field discontinuities, we carefully choose and analyze challenging scenes with significant depth variations from commonly used datasets that exhibit different motion parallax characteristics.
@@
As shown in Fig. 6, the NeRF Synthetic dataset benefits from controlled capture conditions with relatively uniform depth distributions within each scene. In contrast, LLFF and Mip-NeRF 360 datasets contain in-the-wild scenes with larger depth variations. These scenes create sharp phase boundaries where the complex field changes abruptly in between viewpoints, breaking the spatial coherence assumptions of intensity-based representations. While the 3DGS + U-Net method maintains visual quality by first modeling geometry and appearance through intensity-based 3DGS (which handles motion parallax effectively) and then estimating hologram as a secondary task, this two-stage CGH approach sacrifices the scene geometry-awareness across different viewpoints. This limitation highlights the trade-off between the spatial coherence requirements of intensity-based representations and the phase discontinuities introduced by motion parallax in complex-valued holographic radiance fields.
Table 4. Quality comparison on Mip-NeRF 360. 3DGS + U-Net (GT) refers to the Image Quality Ground Truth that requires recalculation of hologram per viewpoint. 3DGS + U-Net (Vary) represents the Viewpoint Relationship Variation without recalculation. Metrics are reported as PSNR/SSIM/LPIPS, respectively.